how do you factor 3x^2-4x+7?
Found 2 solutions by Edwin McCravy, checkley79:
Answer by Edwin McCravy(20056)


how do you factor 3x^2-4x+7
Multiply 3 by 7, getting 21

Think of all the ways to factor 21 using integers:

21 x 1 = 21
 7 x 3 = 21

Since the 7 is positive, list their sums

21 + 1 = 22
 7 + 3 = 10


Since the middle coefficient in absolute value, 4 is
not listed then the trinomial doesn't factor.  So the 
trinomial is prime.

Edwin

Answer by checkley79(3341)

3x^2-4x+7
Using the following quadratic equation
{{x = (-b +- sqrt( b^2-4*a*c ))/(2*a) }}}
x=(4+-sqrt[-4^2-4*3*7])/2*3
x=(4+-sqrt[16-84])/6
x=(4+-sqrt-68)/6
x=(4+-8.246i)/6
x=4/6+-8.246i
x=.666+-1.374i ans.